LOG-IN
Displaying report 1-1 of 1.
VIS (BS)
Print this report.
enzo.tapia - 12:43, Wednesday 28 November 2018 (7162)Get code to link to this report
BS recovery work

As for the beginning of this week, several places in the BS were touching or out of range.

So first I started by trying to move the F0 keystone which is at -2813 um and BF Keystone  which was at -1103 um using the actuation coil but this was not enough.Therefore I adjusted the height using the fishing rod only for the BF, because for now we can not move the F0 FR due to mechanical issues.

Then I tried to move remotely the IM stage in pitch, roll and yaw. Once I finished with pitch and roll. I tried to move the IM in yaw but I could not move it.

Before trying to move yaw with the joystick inside the cleanbooth, I took some TF of the IM and they looked fine. See pictures attached.

After noticing that the adapter cable of the IM picomotor was broken (pins for yaw motion disconnected), So I replaced the adapter and I moved the IM in yaw.

But after that the TF did not look good anymore, and we believe there is something touching again at the IM stage. For example the TF of IM_Y_after. See attached picture.


Moving the IM back in yaw but there was no improvement in the measured TFs. So there are still things to be investigated. And in this context we noticed a clear motion in the IP stage that happened at the same time I moved the IM in yaw. See last attachement.

To be continued...

Images attached to this report
7162_20181128042721_bsiml201811271.png 7162_20181128042732_bsimp201811271.png 7162_20181128042737_bsimr01811271.png 7162_20181128042741_bsimt201811271.png 7162_20181128042745_bsimv201811271.png 7162_20181128042750_bsimy201811271.png 7162_20181128044716_bsimy201811272after.png 7162_20181128044720_bfdof20181127.png
Comments related to this report
enzo.tapia - 14:17, Thursday 29 November 2018 (7190)

I tried to move the horizontal stepper motors of the IP stage in order to center the Preisolator, but the network of the PI steppermotor driver did not work. And it turned out to be that the circuit board was damaged and burned out, probably because of water (maybe from May-June?).

So I borrowed the PI steppermotor driver from SR3 and I have reconfigured the IP address. (10.68.150.56  ----> 10.68.150.52) to keep the same group of address for the BS. The IP address list can be found here: http://gwwiki.icrr.u-tokyo.ac.jp/JGWwiki/KAGRA/Subgroups/DGS/Projects/StepperMotor

I also opened the chassis and checked circuit boards of the GAS steppermotor driver, the TM satellite amplifier and BF picomotor driver of the BS and they are all fine. 

 

 

enzo.tapia - 16:32, Friday 30 November 2018 (7208)

Pictures of some of the circuit boards used at the BS area, second floor.

I inspected 1 Stepper motor driver (PI), 1 Satellite Amp (TM), and 1 picomotor driver (BF cap).

 

The last 2 pictures are from the damaged stepper motor driver.

enzo.tapia - 16:52, Monday 03 December 2018 (7266)

Last Friday I also moved the F1 GAS downwards to ~0 µm using the fishing rod (FR).

Summary: 

  • BF moved upwards from -1180 µm to ~0 µm  using the FR (-470.000 steps).
  • F1 moved downwards from 320 µm to ~0 µm using its FR (550.000 steps). See figure 1.

Then I measured the power spectra of all the sensors (except the oplev QPDs), and the BS seems to be freely swinging except for the F0 GAS, which will tbe fixed when we can open the tank in January. See figure 2.

However, this was not the final height, because the temperature of the BS changed over the weekend from 25.5~26.0 to 24.0ºC, and therefore the BF keystone moved up to 360 µm and the F1 keystone moved up to 400 µm. See figure 3.

And finally I measured again the power spectra of the sensors and the BS is still freely swinging. See figure 4. In fact the temperature change helped to raise the F0 keystone, which is now bouncing free, but still very low (at -2550 µm), the keystone can be raised up by using the coil magnet acutator, but only up to -1800 µm (We can't use the its FR for now). See figure 5.